# Retrieval-Augmented Generation with Estimation of Source Reliability

### v3

@@ Abstract
Retrieval-Augmented Generation (RAG) is an effective approach to enhance the factual accuracy of large language models (LLMs) by retrieving information from external databases, which are typically composed of diverse sources, to supplement the limited internal knowledge of LLMs. However, the standard RAG often risks retrieving incorrect information, as it relies solely on relevance between a query and a document, overlooking the heterogeneous reliability of these sources. To address this issue, we propose Reliability-Aware RAG (RA-RAG), a new multi-source RAG framework that estimates the reliability of sources and leverages this information to prioritize highly reliable and relevant documents, ensuring more robust and accurate response generation. Specifically, RA-RAG first estimates source reliability by cross-checking information across multiple sources. It then retrieves documents from the top-$\kappa$ reliable and relevant sources and aggregates their information using weighted majority voting (WMV), where the selective retrieval ensures scalability while not compromising the performance. Comprehensive experiments show that RA-RAG consistently outperforms baselines in scenarios with heterogeneous source reliability while scaling efficiently as the number of sources increases. Furthermore, we demonstrate the ability of RA-RAG to estimate real-world sources’ reliability, highlighting its practical applicability.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across various tasks Zhao et al. (2023b); Brown et al. (2020). However, they often produce incorrect outputs, particularly when handling up-to-date knowledge that is absent from their internal knowledge Shuster et al. (2021); Zhang et al. (2024); Dhuliawala et al. (2024); Huang et al. (2023); Zhao et al. (2023a). To address this limitation, retrieval-augmented generation (RAG) Guu et al. (2020); Lewis et al. (2020); Asai et al. (2024); Yan et al. (2024) has emerged as a promising approach, leveraging external knowledge from large-scale databases that integrate an extensive set of sources Vu et al. (2024); Kasai et al. (2024).
@@
Although external databases provide valuable information, they also risk retrieving incorrect information from unreliable sources Pan et al. (2023); Chen et al. (2024); Greshake et al. (2023). This vulnerability stems from a fundamental limitation of retrieval, which relies solely on relevance measures between queries and documents Robertson and Walker (1994); Karpukhin et al. (2020); Ni et al. (2022); Izacard et al. (2022), overlooking source reliability heterogeneity. Furthermore, malicious sources can exploit this limitation by crafting highly relevant yet incorrect documents, leading to misleading outputs Zhong et al. (2023); Zou et al. (2024). While existing methods Weller et al. (2024); Xiang et al. (2024); Deng et al. (2024); Pan et al. (2024) attempt to mitigate this issue by refining retrieved documents, they do not address the retrieval problem itself, allowing unreliable sources to dominate the retrieval process.
@@ 3.2 Multi-source RAG with source reliability
We introduce a multi-source RAG framework that explicitly distinguishes between the sources of documents and incorporates source reliability. Let $N$ be the number of distinct sources contributing to the database $\mathcal{D}$. We partition the database as $\mathcal{D}=\bigcup_{i=1}^{N}\mathcal{S}_{i}$, where $\mathcal{S}_{i}$ is the set of documents originating from source $i\in[N]$. Such a partition of dataset $\mathcal{D}$ enables the system to account for the reliability of each document’s source, based on weighted majority voting (WMV). For a given query $q$, let $\hat{y}_{i}=\mathcal{G}(q,\mathcal{R}(q,\mathcal{S}_{i}))$ represent the generated response using retrieved documents exclusively from source $S_{i}$. Once the probability of a retrieved document from source $i$ being correct is estimated as $v_{i}$, and a set of candidate responses $\mathcal{M}$ is obtained from $\hat{y}_{i}$’s, we apply WMV to aggregate the responses as follows:
@@ 4.1 Aggregation process
$f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))$
=$=\begin{cases}\text{IDK}&\text{if }\mathcal{E}(\hat{y}_{i};q,\mathcal{R}(q,\mathcal{S}_{i}))<\tau\\ \hat{y}_{i}&\text{otherwise}\end{cases}\;,$
@@
Additionally, since LLMs often paraphrase responses with equivalent meanings (e.g., “There are 24 hours in a day." vs. “Each day has 24 hours."), we cluster responses in $\mathcal{M}_{\text{filtered}}$ based on semantic equivalence. We denote the refined set obtained through semantic clustering as $\mathcal{C}(\mathcal{M}_{\text{filtered}})=\{C_{k}\subseteq\mathcal{M}_{\text{filtered}}\}_{k=1}^{K}$, where each $C_{k}$ represents a distinct cluster such that $C_{i}\bigcap C_{j}=\emptyset$ for all i $\neq$ j. For semantic clustering method $\mathcal{C}$, we employ the algorithm by Kuhn et al. (2023), which clusters responses that mutually entail each other using a pretrained natural language inference (NLI) model. Following Kuhn et al. (2023), we use the DeBERTa-large model He et al. (2021) for clustering.
@@ 4.2 Iterative reliability estimation
To estimate source reliability and effectively aggregate outputs, we utilize the WMV method proposed by Li and Yu (2014), a simple yet effective approach for aggregating crowdsourced labels in classification tasks. Specifically, we first generate fact-checking queries for documents. For example, if a document in a source states, “COVID-19 is caused by 5G networks", we can generate a query such as “What causes COVID-19?". Given a set of $M$ fact-checking queries, denoted as $\{q^{j}\mid j\in[M]\}$, the iterative reliability estimation process is described as follows:
@@ 5.2 Main results
Figure 3 highlights the importance of considering source reliability when aggregating information across sources. While MV selects “November 2018" based only on response frequency, although it has low reliability, RA-RAG correctly identifies “November 2017" by leveraging well-estimated source reliabilities.
@@ 5.3 Ablation studies and analysis
# Src $\kappa$-RSS $\kappa$-RRSS $5$ $0.588$ $0.597$ $10$ $0.663$ $0.727$ $1000$ $0.689$ $0.768$
@@ 6 Real-world Application
# Src $\kappa$-RRSS Token Consumption ($\downarrow$) API Calls ($\downarrow$) Inference Cost ($\downarrow$) Accuracy ($\uparrow$) $5$ w/o $3138$ $5$ $0.00048$ $0.597$ w/ $3055$ ($\downarrow 2.6\%$) $4.87$ ($\downarrow 2.6\%$) $0.00046$ ($\downarrow 4.2\%$) $0.597$ $10$ w/o $6272$ $10$ $0.00096$ $0.744$ w/ $4251$ ($\downarrow 32.3\%$) $6.79$ ($\downarrow 32.1\%$) $0.00065$ ($\downarrow 32.3\%$) $0.727$ $1000$ w/o $627115$ $1000$ $0.096$ $0.780$ w/ $5415$ ($\downarrow 99.1\%$) $8.66$ ($\downarrow 99.1\%$) $0.00083$ ($\downarrow 99.1\%$) $0.768$
Table 2: Comparison of computational efficiency with and without $\kappa$-RRSS evaluated on the NQ dataset using GPT-4o-mini. The reported values represent the average per query, with the values in parentheses ()($\cdot$)( ⋅ ) indicating the reduction rate achieved with $\kappa$-RRSS.
@@ 6.1 Setup
Reliability estimation process. To apply our reliability estimation method, we generate yes/no fact-checking queries for each collected claim (e.g., “Is it true that {claim}?"), allowing for straightforward cross-checking of claims across multiple sources. We use Google News as the retriever, which provides rich sources for retrieving relevant documents by selecting the top 20 results. GPT-4o-mini is used as the language model to generate responses.
@@ 6.2 Experimental results
Figure 5 further illustrates that our estimated reliability for User A closely matches the oracle reliability across different reliability levels. The PCC of $0.991$ and SRCC of $0.992$ (both with p-values $<0.001$) indicate a strong correlation. Additionally, an average accuracy of $0.801$ demonstrates the effectiveness of our method in validating claims across varying reliability levels.
@@ 7 Conclusion
8 Limitations
@@ Appendix A Details of Misalignment Filtration
=$=\begin{cases}\text{IDK}&\text{if }\mathcal{E}(\hat{y}_{i};q,\mathcal{R}(q,\mathcal{S}_{i}))<\tau\\ \hat{y}_{i}&\text{otherwise}\end{cases}\;,$
@@ Appendix B $\kappa$-Reliable and Relevant Source Selection ($\kappa$-RRSS)
2:  $\mathcal{M}_{\kappa}\leftarrow\emptyset$
3:  $\text{count}\leftarrow 0$
@@
Types of Retrieved Documents Types of Answers Filtering ($f_{\text{align}}$) Factual Misinformation Irrelevant Correct w/o $96.38$ $5.05$ $26.07$ w/ $94.32$ $2.53$ ($-2.52$) $4.16$ ($-21.93$) Incorrect w/o -- $75.76$- w/ -- $70.96$- IDK w/o $0.26$ $4.80$ $50.92$ w/ $2.58$ $13.89$ ($+9.09$) $91.19$ ($+40.27$) Hallucination w/o $8.01$ $10.10$ $22.89$ w/ $7.75$ $8.33$ ($-1.77$) $4.53$ ($-18.36$)
@@ I.2 Prompt for misinformation generation
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o -- 80.36 2.02 4.53 0.1 76.74 0.76 1.84 w/ 0.5 72.61 0.51 1.35 0.8 68.22 0.25 0.98 Incorrect w/o -- -- 87.63 -- 0.1 -- 86.87 -- w/ 0.5 -- 84.60 -- 0.8 -- 82.32 -- IDK w/o -- 2.33 1.01 80.29 0.1 6.20 4.55 91.31 w/ 0.5 11.89 7.83 94.25 0.8 17.57 10.86 94.86 Hallucination w/o -- 17.31 9.34 15.18 0.1 17.05 7.83 6.85 w/ 0.5 15.50 7.07 4.41 0.8 14.21 6.57 4.16
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o -- 96.38 5.05 26.07 0.1 94.32 2.53 4.16 w/ 0.5 89.41 1.26 1.71 0.8 84.50 1.01 0.73 Incorrect w/o -- -- 75.76 -- 0.1 -- 70.96 -- w/ 0.5 -- 66.67 -- 0.8 -- 62.12 -- IDK w/o -- 0.26 4.80 50.92 0.1 2.58 13.89 91.19 w/ 0.5 8.01 20.20 96.57 0.8 13.44 25.76 98.04 Hallucination w/o -- 8.01 10.10 22.89 0.1 7.75 8.33 4.53 w/ 0.5 7.24 7.58 1.59 0.8 6.72 6.82 1.10
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o -- 82.95 6.57 14.69 0.1 75.97 3.79 2.20 w/ 0.5 66.93 2.27 0.73 0.8 58.66 1.01 0.24 Incorrect w/o -- -- 65.40 -- 0.1 -- 54.55 -- w/ 0.5 -- 45.45 -- 0.8 -- 35.61 -- IDK w/o -- 0.26 8.59 59.24 0.1 9.30 26.01 91.43 w/ 0.5 20.93 41.16 96.94 0.8 31.01 55.05 98.16 Hallucination w/o -- 17.57 16.16 27.29 0.1 15.50 12.37 7.59 w/ 0.5 12.92 7.83 3.55 0.8 11.11 5.05 2.82
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o -- 80.36 2.02 4.53 0.1 78.81 0.25 2.20 w/ 0.5 72.61 0.51 1.35 0.8 68.22 0.25 0.98 Incorrect w/o -- -- 87.63 -- 0.1 -- 89.65 -- w/ 0.5 -- 84.60 -- 0.8 -- 82.32 -- IDK w/o -- 2.33 1.01 80.29 0.1 6.20 3.03 88.13 w/ 0.5 11.89 7.83 94.25 0.8 17.57 10.86 94.86 Hallucination w/o -- 17.31 9.34 15.18 0.1 14.99 7.07 9.67 w/ 0.5 15.50 7.07 4.41 0.8 14.21 6.57 4.16
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o -- 96.38 4.80 36.72 0.1 93.80 1.77 5.51 w/ 0.5 89.41 1.26 1.71 0.8 84.50 1.01 0.73 Incorrect w/o -- -- 77.27 -- 0.1 -- 72.98 -- w/ 0.5 -- 66.67 -- 0.8 -- 62.12 -- IDK w/o -- 0.26 4.55 32.56 0.1 2.84 13.38 89.47 w/ 0.5 8.01 20.20 96.57 0.8 13.44 25.76 98.04 Hallucination w/o -- 8.01 9.09 30.60 0.1 8.01 7.58 4.90 w/ 0.5 7.24 7.58 1.59 0.8 6.72 6.82 1.10
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o -- 83.98 4.55 18.73 0.1 77.00 3.54 3.67 w/ 0.5 68.22 2.53 0.86 0.8 58.91 1.26 0.49 Incorrect w/o -- -- 74.24 -- 0.1 -- 61.62 -- w/ 0.5 -- 48.74 -- 0.8 -- 37.63 -- IDK w/o -- 1.03 3.79 39.53 0.1 10.34 22.47 87.39 w/ 0.5 21.96 40.91 96.08 0.8 32.82 55.30 97.92 Hallucination w/o -- 15.76 14.14 42.96 0.1 13.44 9.09 10.16 w/ 0.5 10.59 4.55 4.28 0.8 9.04 2.53 2.82
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o -- 78.55 0.51 1.84 0.1 75.45 0.51 0.86 w/ 0.5 71.58 0.51 0.61 0.8 67.70 0.00 0.49 Incorrect w/o -- -- 88.64 -- 0.1 -- 87.88 -- w/ 0.5 -- 85.61 -- 0.8 -- 83.08 -- IDK w/o -- 4.91 4.04 92.78 0.1 8.27 5.05 95.10 w/ 0.5 13.18 7.58 95.84 0.8 18.09 10.86 95.96 Hallucination w/o -- 16.54 6.82 5.39 0.1 16.28 6.57 4.04 w/ 0.5 15.25 6.31 3.55 0.8 14.21 6.06 3.55
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o -- 96.64 1.52 9.67 0.1 94.32 0.76 2.45 w/ 0.5 89.41 0.51 1.10 0.8 84.75 0.25 0.86 Incorrect w/o -- -- 68.43 -- 0.1 -- 66.41 -- w/ 0.5 -- 62.63 -- 0.8 -- 59.09 -- IDK w/o -- 0.78 16.92 87.76 0.1 3.10 21.21 96.82 w/ 0.5 8.53 25.76 98.41 0.8 13.70 30.30 98.65 Hallucination w/o -- 7.24 8.84 2.45 0.1 7.24 7.32 0.61 w/ 0.5 6.72 6.82 0.37 0.8 6.20 6.06 0.37
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o -- 86.30 5.56 9.42 0.1 78.81 3.03 1.71 w/ 0.5 69.51 2.27 0.73 0.8 59.69 1.01 0.37 Incorrect w/o -- -- 63.38 -- 0.1 -- 54.29 -- w/ 0.5 -- 45.20 -- 0.8 -- 36.11 -- IDK w/o -- 0.78 16.67 85.19 0.1 9.30 30.30 96.21 w/ 0.5 20.93 42.93 97.92 0.8 31.78 55.56 98.41 Hallucination w/o -- 13.70 11.11 6.61 0.1 12.66 9.09 3.30 w/ 0.5 10.34 6.31 2.57 0.8 9.30 4.04 2.45

### v4

@@ Abstract
Retrieval-Augmented Generation (RAG) is an effective approach to enhance the factual accuracy of large language models (LLMs) by retrieving information from external databases, which are typically composed of diverse sources, to supplement the limited internal knowledge of LLMs. However, the standard RAG often risks retrieving incorrect information, as it relies solely on relevance between a query and a document, overlooking the heterogeneous reliability of these sources. To address this issue, we propose Reliability-Aware RAG (RA-RAG), a new multi-source RAG framework that estimates the reliability of sources and leverages this information to prioritize highly reliable and relevant documents, ensuring more robust and accurate response generation. Specifically, RA-RAG first estimates source reliability by cross-checking information across multiple sources. It then retrieves documents from the top-$\kappa$ reliable and relevant sources and aggregates their information using weighted majority voting (WMV), where the selective retrieval ensures scalability while not compromising the performance. Comprehensive experiments show that RA-RAG consistently outperforms baselines in scenarios with heterogeneous source reliability while scaling efficiently as the number of sources increases. Furthermore, we demonstrate the ability of RA-RAG to estimate real-world sources’ reliability, highlighting its practical applicability. Our code and data are available at RA-RAG.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across various tasks Zhao et al. (2023b); Brown et al. (2020). However, they often produce incorrect outputs, particularly when handling up-to-date knowledge that is absent from their internal knowledge Shuster et al. (2021); Zhang et al. (2024); Dhuliawala et al. (2024); Huang et al. (2023); Zhao et al. (2023a). To address this limitation, retrieval-augmented generation (RAG) Guu et al. (2020); Lewis et al. (2020); Asai et al. (2024); Yan et al. (2024) has emerged as a promising approach, leveraging external knowledge from large-scale databases that integrate an extensive set of sources to enhance coverage and enable richer responses Vu et al. (2024); Kasai et al. (2024).
@@
However, while such databases provide valuable information, they also risk retrieving incorrect information from unreliable sources Pan et al. (2023); Chen et al. (2024); Greshake et al. (2023). Moreover, even PerplexityPerplexity AI (2025), state-of-the-art commercial RAG systems, have been observed to spread misinformation by retrieving content from AI-generated spam blogs Shrivastava (2024).
This vulnerability stems from a fundamental limitation of retrieval, which relies solely on relevance measures between queries and documents Robertson and Walker (1994); Karpukhin et al. (2020); Ni et al. (2022); Izacard et al. (2022), overlooking source reliability heterogeneity. Furthermore, malicious sources can exploit this limitation by crafting highly relevant yet incorrect documents, leading to misleading outputs Zhong et al. (2023); Zou et al. (2024). While existing methods Weller et al. (2024); Xiang et al. (2024); Deng et al. (2024); Pan et al. (2024) attempt to mitigate this issue by refining retrieved documents, they do not address the retrieval problem itself, allowing unreliable sources to dominate the retrieval process.
@@ 3.2 Multi-source RAG with source reliability
We introduce a multi-source RAG framework that explicitly distinguishes between the sources of documents and incorporates source reliability. Let $N$ be the number of distinct sources contributing to the database $\mathcal{D}$. We partition the database as $\mathcal{D}=\bigcup_{i=1}^{N}\mathcal{S}_{i}$, where $\mathcal{S}_{i}$ is the set of documents originating from source $i\in[N]$. The definition of a “source” is application-dependent and may vary in granularity: sources can be fine-grained (e.g., individual social media accounts or statements by specific individuals such as politicians) or coarse-grained (e.g., news websites). In Section 6, we demonstrate practical applications of this framework.
This partitioning enables the system to account for the reliability of each document’s source, based on weighted majority voting (WMV). For a given query $q$, let $\hat{y}_{i}=\mathcal{G}(q,\mathcal{R}(q,\mathcal{S}_{i}))$ represent the generated response using retrieved documents exclusively from source $S_{i}$. Once the probability of a retrieved document from source $i$ being correct is estimated as $v_{i}$, and a set of candidate responses $\mathcal{M}$ is obtained from $\hat{y}_{i}$’s, we apply WMV to aggregate the responses as follows:
@@ 4.1 Aggregation process
$f_{\text{align}}(\hat{y}_{i},q,\mathcal{R}(q,\mathcal{S}_{i}))$ (2)
[…]
@@
Additionally, since LLMs often paraphrase responses with equivalent meanings (e.g., “There are 24 hours in a day." vs. “Each day has 24 hours.”), we cluster responses in $\mathcal{M}_{\text{filtered}}$ based on semantic equivalence. We denote the refined set obtained through semantic clustering as $\mathcal{C}(\mathcal{M}_{\text{filtered}})=\{C_{k}\subseteq\mathcal{M}_{\text{filtered}}\}_{k=1}^{K}$, where each $C_{k}$ represents a distinct cluster such that $C_{i}\bigcap C_{j}=\emptyset$ for all i $\neq$ j. For semantic clustering method $\mathcal{C}$, we employ the algorithm by Kuhn et al. (2023), which clusters responses that mutually entail each other using a pretrained natural language inference (NLI) model. Following Kuhn et al. (2023), we use the DeBERTa-large model He et al. (2021) for clustering.
@@ 4.2 Iterative reliability estimation
To estimate source reliability and effectively aggregate outputs, we extend the WMV method proposed by Li and Yu (2014), a simple yet effective approach for aggregating crowdsourced labels in classification tasks. Specifically, we first generate fact-checking queries for documents. For example, if a document in a source states, “COVID-19 is caused by 5G networks”, we can generate a query such as “What causes COVID-19?”. Given a set of $M$ fact-checking queries, denoted as $\{q^{j}\mid j\in[M]\}$, the iterative reliability estimation process is described as follows:
@@ 5.2 Main results
Figure 3 highlights the importance of considering source reliability when aggregating information across sources. While MV selects “November 2018” based only on response frequency, although it has low reliability, RA-RAG correctly identifies “November 2017” by leveraging well-estimated source reliabilities.
@@ 5.3 Ablation studies and analysis
# Src $\kappa$-RSS $\kappa$-RRSS 55 $0.588$ $0.597$ $10$ $0.663$ $0.727$ $1000$ $0.689$ $0.768$
@@ 6 Real-world Application
# Src $\kappa$-RRSS Token Consumption ($\downarrow$) API Calls ($\downarrow$) Inference Cost ($\downarrow$) Accuracy ($\uparrow$) 55 w/o $3138$ 55 $0.00048$ $0.597$ w/ $3055$ ($\downarrow 2.6\%$) $4.87$ ($\downarrow 2.6\%$) $0.00046$ ($\downarrow 4.2\%$) $0.597$ $10$ w/o $6272$ $10$ $0.00096$ $0.744$ w/ $4251$ ($\downarrow 32.3\%$) $6.79$ ($\downarrow 32.1\%$) $0.00065$ ($\downarrow 32.3\%$) $0.727$ $1000$ w/o $627115$ $1000$ $0.096$ $0.780$ w/ $5415$ ($\downarrow 99.1\%$) $8.66$ ($\downarrow 99.1\%$) $0.00083$ ($\downarrow 99.1\%$) $0.768$
Table 2: Comparison of computational efficiency with and without $\kappa$-RRSS evaluated on the NQ dataset using GPT-4o-mini. The reported values represent the average per query, with the values in parentheses $(\cdot)$ indicating the reduction rate achieved with $\kappa$-RRSS.
@@ 6.1 Setup
Reliability estimation process. To apply our reliability estimation method, we generate yes/no fact-checking queries for each collected claim (e.g., “Is it true that {claim}?”), allowing for straightforward cross-checking of claims across multiple sources. We use Google News as the retriever, which provides rich sources for retrieving relevant documents by selecting the top 20 results. GPT-4o-mini is used as the language model to generate responses.
@@ 6.2 Experimental results
Figure 5 further illustrates that our estimated reliability for User A closely matches the oracle reliability across different reliability levels. The PCC of $0.991$ and SRCC of $0.992$ (both with p-values <0.001<0.001) indicate a strong correlation. Additionally, an average accuracy of $0.801$ demonstrates the effectiveness of our method in validating claims across varying reliability levels.
@@ 7 Conclusion
While our work focuses on short-form question answering, the RA-RAG framework extends to more complex tasks like long-form and multi-hop question answering. For example, in biography generation, an LLM can first decompose the task into atomic subquestions (e.g., “What is the person’s age?”; “Where did the person live?”). Then, our framework can be applied to obtain reliable answers to each of these short questions, after which an LLM aggregates these answers to compose a coherent long-form response. A promising direction for future work is the construction of datasets that comprise diverse sources with heterogeneous reliability for long-form and multi-hop question answering, enabling more rigorous evaluation and the development of stronger source-aware method
Limitations
@@
Acknowledgments
This work was supported by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (No. RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH)); the IITP grant funded by the Korean government (MSIT) (No. RS-2024-00509258, Global AI Frontier Lab); the IITP–ITRC (Information Technology Research Center) grant funded by the Korean government (MSIT) (No. IITP-2025-00437866); the IITP grant funded by the Korean government (MSIT) (No. RS-2024-00457882, AI Research Hub Project); the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. RS-2023-00217286); and the NRF grant funded by the Korean government (MSIT) (No. RS-2025-00560062).
@@ Appendix A Details of Misalignment Filtration
$=\begin{cases}\text{IDK}&\text{if }\mathcal{E}(\hat{y}_{i};q,\mathcal{R}(q,\mathcal{S}_{i}))<\tau\\ \hat{y}_{i}&\text{otherwise}\end{cases}\;,$
@@ Appendix B $\kappa$-Reliable and Relevant Source Selection ($\kappa$-RRSS)
[…]
@@
Types of Retrieved Documents Types of Answers Filtering ($f_{\text{align}}$) Factual Misinformation Irrelevant Correct w/o $96.38$ $5.05$ $26.07$ w/ $94.32$ $2.53$ (2.52-2.52) $4.16$ (21.93-21.93) Incorrect w/o - $75.76$ - w/ - $70.96$ - IDK w/o $0.26$ $4.80$ $50.92$ w/ $2.58$ $13.89$ (+9.09+9.09) $91.19$ (+40.27+40.27) Hallucination w/o $8.01$ $10.10$ $22.89$ w/ $7.75$ $8.33$ (1.77-1.77) $4.53$ (18.36-18.36)
@@ I.2 Prompt for misinformation generation
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o - 80.36 2.02 4.53 0.1 76.74 0.76 1.84 w/ 0.5 72.61 0.51 1.35 0.8 68.22 0.25 0.98 Incorrect w/o - - 87.63 - 0.1 - 86.87 - w/ 0.5 - 84.60 - 0.8 - 82.32 - IDK w/o - 2.33 1.01 80.29 0.1 6.20 4.55 91.31 w/ 0.5 11.89 7.83 94.25 0.8 17.57 10.86 94.86 Hallucination w/o - 17.31 9.34 15.18 0.1 17.05 7.83 6.85 w/ 0.5 15.50 7.07 4.41 0.8 14.21 6.57 4.16
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o - 96.38 5.05 26.07 0.1 94.32 2.53 4.16 w/ 0.5 89.41 1.26 1.71 0.8 84.50 1.01 0.73 Incorrect w/o - - 75.76 - 0.1 - 70.96 - w/ 0.5 - 66.67 - 0.8 - 62.12 - IDK w/o - 0.26 4.80 50.92 0.1 2.58 13.89 91.19 w/ 0.5 8.01 20.20 96.57 0.8 13.44 25.76 98.04 Hallucination w/o - 8.01 10.10 22.89 0.1 7.75 8.33 4.53 w/ 0.5 7.24 7.58 1.59 0.8 6.72 6.82 1.10
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o - 82.95 6.57 14.69 0.1 75.97 3.79 2.20 w/ 0.5 66.93 2.27 0.73 0.8 58.66 1.01 0.24 Incorrect w/o - - 65.40 - 0.1 - 54.55 - w/ 0.5 - 45.45 - 0.8 - 35.61 - IDK w/o - 0.26 8.59 59.24 0.1 9.30 26.01 91.43 w/ 0.5 20.93 41.16 96.94 0.8 31.01 55.05 98.16 Hallucination w/o - 17.57 16.16 27.29 0.1 15.50 12.37 7.59 w/ 0.5 12.92 7.83 3.55 0.8 11.11 5.05 2.82
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o - 80.36 2.02 4.53 0.1 78.81 0.25 2.20 w/ 0.5 72.61 0.51 1.35 0.8 68.22 0.25 0.98 Incorrect w/o - - 87.63 - 0.1 - 89.65 - w/ 0.5 - 84.60 - 0.8 - 82.32 - IDK w/o - 2.33 1.01 80.29 0.1 6.20 3.03 88.13 w/ 0.5 11.89 7.83 94.25 0.8 17.57 10.86 94.86 Hallucination w/o - 17.31 9.34 15.18 0.1 14.99 7.07 9.67 w/ 0.5 15.50 7.07 4.41 0.8 14.21 6.57 4.16
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o - 96.38 4.80 36.72 0.1 93.80 1.77 5.51 w/ 0.5 89.41 1.26 1.71 0.8 84.50 1.01 0.73 Incorrect w/o - - 77.27 - 0.1 - 72.98 - w/ 0.5 - 66.67 - 0.8 - 62.12 - IDK w/o - 0.26 4.55 32.56 0.1 2.84 13.38 89.47 w/ 0.5 8.01 20.20 96.57 0.8 13.44 25.76 98.04 Hallucination w/o - 8.01 9.09 30.60 0.1 8.01 7.58 4.90 w/ 0.5 7.24 7.58 1.59 0.8 6.72 6.82 1.10
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o - 83.98 4.55 18.73 0.1 77.00 3.54 3.67 w/ 0.5 68.22 2.53 0.86 0.8 58.91 1.26 0.49 Incorrect w/o - - 74.24 - 0.1 - 61.62 - w/ 0.5 - 48.74 - 0.8 - 37.63 - IDK w/o - 1.03 3.79 39.53 0.1 10.34 22.47 87.39 w/ 0.5 21.96 40.91 96.08 0.8 32.82 55.30 97.92 Hallucination w/o - 15.76 14.14 42.96 0.1 13.44 9.09 10.16 w/ 0.5 10.59 4.55 4.28 0.8 9.04 2.53 2.82
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o - 78.55 0.51 1.84 0.1 75.45 0.51 0.86 w/ 0.5 71.58 0.51 0.61 0.8 67.70 0.00 0.49 Incorrect w/o - - 88.64 - 0.1 - 87.88 - w/ 0.5 - 85.61 - 0.8 - 83.08 - IDK w/o - 4.91 4.04 92.78 0.1 8.27 5.05 95.10 w/ 0.5 13.18 7.58 95.84 0.8 18.09 10.86 95.96 Hallucination w/o - 16.54 6.82 5.39 0.1 16.28 6.57 4.04 w/ 0.5 15.25 6.31 3.55 0.8 14.21 6.06 3.55
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o - 96.64 1.52 9.67 0.1 94.32 0.76 2.45 w/ 0.5 89.41 0.51 1.10 0.8 84.75 0.25 0.86 Incorrect w/o - - 68.43 - 0.1 - 66.41 - w/ 0.5 - 62.63 - 0.8 - 59.09 - IDK w/o - 0.78 16.92 87.76 0.1 3.10 21.21 96.82 w/ 0.5 8.53 25.76 98.41 0.8 13.70 30.30 98.65 Hallucination w/o - 7.24 8.84 2.45 0.1 7.24 7.32 0.61 w/ 0.5 6.72 6.82 0.37 0.8 6.20 6.06 0.37
@@
Types of Retrieved Documents Types of Answers Filtering $(f_{\text{align}})$ Threshold Factual Misinformation Irrelevant Correct w/o - 86.30 5.56 9.42 0.1 78.81 3.03 1.71 w/ 0.5 69.51 2.27 0.73 0.8 59.69 1.01 0.37 Incorrect w/o - - 63.38 - 0.1 - 54.29 - w/ 0.5 - 45.20 - 0.8 - 36.11 - IDK w/o - 0.78 16.67 85.19 0.1 9.30 30.30 96.21 w/ 0.5 20.93 42.93 97.92 0.8 31.78 55.56 98.41 Hallucination w/o - 13.70 11.11 6.61 0.1 12.66 9.09 3.30 w/ 0.5 10.34 6.31 2.57 0.8 9.30 4.04 2.45